# Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values

### v2

@@ Appendix C Quantization and Dequantization Error Bounds
This result implies that, for a given bit-width $b$, matrices with larger spectral norms incur proportionally larger quantization errors. Consequently, a matrix that exhibits a larger spectral norm is more susceptible to quantization errors. To control the error propagation in such matrices, a higher bit-width is necessary.
@@
C.1 Preliminaries and Notation
Let $A\in\mathbb{R}^{m\times n}$ denote a real matrix whose entries are to be stored using a fixed number of bits in two’s-complement representation. The following notation is adopted:
@@ C.1 Preliminaries and Notation
    $Q_{ij}$ $=\mathrm{round}\!\Bigl{(}\frac{A_{ij}}{\alpha}\Bigr{)},$ (18)
@@
C.2 Spectral Norm Bound
Theorem 3 (Spectral Norm Quantization Error).
Let $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{N}$ be given. Consider two’s-complement quantization with scale factor
$\alpha\;=\;\frac{M}{2^{b-1}-1},\quad M\;=\;\max_{i,j}|A_{ij}|.$
Define
$Q_{ij}\;=\;\mathrm{round}\!\Bigl{(}\frac{A_{ij}}{\alpha}\Bigr{)}\quad\text{and}\quad\widehat{A}_{ij}\;=\;\alpha\,Q_{ij}.$
Then, the following bound holds:
$M$ $\leq\|A-\widehat{A}\|_{2}\leq\sqrt{m\,n}\,\frac{M}{2(2^{b-1}-1)}$ (19)
$\leq\sqrt{m\,n}\,\frac{\|A\|_{2}}{2(2^{b-1}-1)}.$
In approximate form for large $b$,
$\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
@@ C.2 Spectral Norm Bound
C.3 Frobenius Norm Bound
Theorem 4 (Frobenius Norm Quantization Error).
Under the same setup, the Frobenius norm of the quantization error satisfies
$\begin{split}\|A-\widehat{A}\|_{F}&\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{M}{2(2^{b-1}-1)}}\\ &\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{\|A\|_{F}}{2(2^{b-1}-1)}}.\end{split}$ (20)
In approximate form,
$\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
@@ C.3 Frobenius Norm Bound
$\|A-\widehat{A}\|_{2}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2},$ (21)
@@ C.4 Quantizing KV Cache
C.5 Spectral-Norm Perspective
Standard quantization error bounds yield
$\|\mathbf{E}_{V}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{2},$ (22)
@@ C.5 Spectral-Norm Perspective
C.6 Frobenius-Norm (MSE) Perspective
Similarly, the Frobenius norm error bounds are given by
$\|\mathbf{E}_{V}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F},$ (23)
$\|\mathbf{E}_{K}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F}.$
To achieve comparable Frobenius errors,
@@ C.6 Frobenius-Norm (MSE) Perspective
which implies
[…]
and therefore,
@@
Again, because $\|\mathbf{V}\|_{F}<\|\mathbf{K}\|_{F}$, it follows that $b_{K}>b_{V}$.
Appendix D Supplementary Key-Value Weight Norm Analysis
Here we present additional evidence supporting the Key-Value Norm Disparity theorem established in Section 3.1. While Fig. 2 demonstrates this observation for the Llama model family, the same pattern consistently emerges across Mistral family.
@@ Appendix D Supplementary Key-Value Weight Norm Analysis
Figure 4: Frobenius norm plot for Mistral Family. The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude. The spectral norms are higher for the key weights than for the value weights across layers.
Appendix E Extended Visualization of Singular Value Distributions
@@
Figure 5 provides the full-range view of the singular value distribution for the key and value cache at various layers of the Llama 3.3 70B model on the C4 dataset. The horizontal axis enumerates the singular values in descending order, beginning with the largest (i.e., the spectral norm), and the vertical axis measures the corresponding magnitudes. The shaded region captures the minimum-to-maximum range of singular values across the attention heads within each layer, while the solid curves depict the mean singular value at each rank.
@@
Figure 5: Complete singular value distribution of key and value cache for the Llama 3.3 70B model on the C4 dataset. The x-axis denotes the singular value indices, ordered from the largest (spectral norm) to the smallest, while the y-axis represents the corresponding magnitudes. The shaded region illustrates the range between the minimum and maximum singular values across attention heads within each layer, and the solid lines indicate the mean singular value magnitude at each index. This full-spectrum view highlights that key matrices consistently maintain significantly higher singular values throughout the entire distribution, further reinforcing their dominant representational capacity compared to value matrices.
@@
Figure 6: Frobenius norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the layer index. The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer. Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
Figure 7: Spectral norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the layer index. The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer. Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
Appendix F Extended Quantization Error Results
@@
Table 4: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the MMLU dataset.
Model K2 V2 K3 V3 K4 V4
@@
Table 5: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the C4 dataset.
Model K2 V2 K3 V3 K4 V4
@@
Model K2 V2 K3 V3 K4 V4
@@
Figure 8: Mean squared error (MSE) of KV quantization for Llama 3.3-70B on C4. We use quantization bit-widths ranging from 2 to 8. The x-axis represents the quantization bit-width, while the y-axis shows the MSE on a logarithmic scale. A logarithmic scale is used to highlight differences at higher bit-widths, where MSE values decrease significantly and approach zero, particularly at 8-bit. Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
@@
(a) C4 dataset
(b) MMLU dataset
(c) GSM8k dataset
Appendix G Reproducibility and Resources
Inference was performed using the Hugging Face Transformers [53] and Accelerate [54] libraries, together with FlashAttention [55, 56], to optimize token-generation throughput. We integrated both Quanto and HQQ into the Language Model Evaluation Harness [57] to enable systematic and reproducible evaluation of model performance under varying quantization schemes. All runs were executed on GPU clusters equipped with NVIDIA A100 (80 GB) using the Quanto backend and NVIDIA H200 (141 GB) using the HQQ backend.
Appendix H Extended Evaluation Results
@@
Extended evaluation results for KV cache quantization under the HQQ backend are presented in Tables 910 (GSM8K flexible and strict accuracies), Tables 8 (EQ-Parseable parseable-pass rates and exact-match scores), and Table 7 (CoQA F1 scores).
Table 7: Full CoQA Accuracies F1.
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0.129657 0.147672 0.127875 0.222218 0.127204 0.358652 0.272695 0.231077
k1v2 0.226803 0.236351 0.222816 0.360911 0.206763 0.384316 0.329965 0.327598
k1v4 0.226441 0.236751 0.229168 0.375805 0.241965 0.385504 0.323921 0.34914
k1v6 0.214566 0.242716 0.229515 0.378638 0.220753 0.378799 0.34787 0.351566
k1v8 0.219723 0.240167 0.221966 0.364095 0.241671 0.382309 0.341797 0.3545
k2v1 0.252746 0.189839 0.576534 0.463689 0.565141 0.598736 0.73744 0.732328
k2v2 0.334262 0.525527 0.760021 0.72138 0.757424 0.766938 0.801972 0.808949
k2v4 0.333106 0.558613 0.765639 0.732337 0.762769 0.760775 0.799323 0.808594
k2v6 0.339966 0.564876 0.764105 0.730226 0.763797 0.765849 0.797442 0.80742
k2v8 0.331698 0.567622 0.763972 0.733084 0.765828 0.769881 0.79908 0.803599
k4v1 0.496626 0.575004 0.768665 0.803298 0.773274 0.815164 0.823299 0.817609
k4v2 0.666081 0.693266 0.797161 0.805846 0.785901 0.822081 0.819957 0.819761
k4v4 0.691803 0.701052 0.797872 0.806737 0.788327 0.818848 0.816583 0.824813
k4v6 0.688022 0.701648 0.797325 0.809305 0.784273 0.817312 0.818533 0.823784
k4v8 0.686547 0.700206 0.79702 0.807328 0.782431 0.817872 0.818202 0.822701
k6v1 0.649633 0.597133 0.770615 0.801156 0.765218 0.812864 0.817002 0.816527
k6v2 0.692575 0.699732 0.798369 0.80343 0.784922 0.820546 0.821854 0.825656
k6v4 0.701547 0.705158 0.79613 0.806657 0.786 0.820736 0.822193 0.824622
k6v6 0.700741 0.705549 0.795475 0.807525 0.790379 0.81943 0.821026 0.825455
k6v8 0.693932 0.702742 0.79642 0.806614 0.789155 0.81943 0.820462 0.825005
k8v1 0.650554 0.598797 0.768704 0.802301 0.767131 0.814654 0.818909 0.818461
k8v2 0.691099 0.696408 0.798332 0.802632 0.787855 0.822552 0.822365 0.824697
k8v4 0.70116 0.703779 0.794958 0.807328 0.788884 0.820734 0.822508 0.825976
k8v6 0.698706 0.70461 0.795362 0.808917 0.786832 0.818963 0.820002 0.825695
k8v8 0.696944 0.70377 0.794663 0.808457 0.785888 0.820653 0.819874 0.826971
bf16 0.707878 0.70117 0.794959 0.807871 0.785389 0.820364 0.81985 0.826088
Table 8: Full EQ-Bench Parseable.
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-32B
k1v1 0 0 0 0 0 0 0
k1v2 0 0 0 0 0 0 0
k1v4 0 0 0 0 0 0 0
k1v6 0 0 0 0 0 0 0
k1v8 0 0 0 0 0 0 0
k2v1 0 0 8.187 0 35.673 0 2.339
k2v2 0 16.374 80.702 27.485 85.965 65.497 67.836
k2v4 0 34.503 84.795 39.766 88.889 70.175 70.175
k2v6 0 39.181 83.041 36.842 90.059 67.836 69.591
k2v8 0 38.012 84.211 40.936 88.889 68.421 70.760
k4v1 0 47.953 58.480 77.778 83.626 98.830 77.778
k4v2 63.743 97.076 95.322 87.719 97.661 95.322 80.702
k4v4 65.497 97.076 98.830 84.795 99.415 95.322 80.702
k4v6 64.328 97.076 98.830 85.380 98.830 93.567 80.702
k4v8 62.573 97.076 98.830 84.795 98.830 93.567 80.702
k6v1 23.392 60.234 53.801 81.287 90.643 96.491 79.532
k6v2 99.415 97.076 94.737 87.719 97.661 94.737 80.702
k6v4 99.415 97.661 99.415 87.135 98.830 95.906 80.702
k6v6 99.415 97.661 98.830 86.550 98.830 96.491 80.702
k6v8 99.415 97.661 98.830 87.135 98.830 97.661 80.702
k8v1 26.901 60.234 56.725 78.947 92.983 97.076 77.193
k8v2 99.415 97.076 94.737 87.135 97.661 96.491 79.532
k8v4 99.415 97.661 99.415 87.719 98.830 95.322 80.702
k8v6 99.415 97.661 99.415 86.550 98.830 95.906 80.702
k8v8 99.415 97.661 99.415 86.550 98.830 95.906 80.702

### v3

@@ Appendix B Quantization Error Bounds
This result implies that, for a given bit-width $b$, matrices with larger spectral norms incur proportionally larger quantization errors. Consequently, a matrix that exhibits a larger spectral norm is more susceptible to quantization errors. To control error propagation in such matrices, a higher bit width is necessary.
@@
B.1 Preliminaries
Let $A\in\mathbb{R}^{m\times n}$ denote a real matrix whose entries are to be stored using a fixed number of bits in two’s complement representation. The following notation is adopted:
@@ B.1 Preliminaries
    $Q_{ij}$ $=\mathrm{round}\!\Bigl(\frac{A_{ij}}{\alpha}\Bigr),$
@@
B.2 Spectral Norm Error Bound
Theorem 3 (Spectral Norm Error Bound for Uniform Quantization) Let $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{N}$ be given. Consider two’s complement quantization with a scale factor of $\alpha\;=\;\frac{M}{2^{b-1}-1},\quad M\;=\;\max_{i,j}|A_{ij}|.$ Define $Q_{ij}\;=\;\mathrm{round}\!\Bigl(\frac{A_{ij}}{\alpha}\Bigr)\quad\text{and}\quad\widehat{A}_{ij}\;=\;\alpha\,Q_{ij}.$ Then, the following bound holds: $M$ $\leq\|A-\widehat{A}\|_{2}\leq\sqrt{m\,n}\,\frac{M}{2(2^{b-1}-1)}$ $\leq\sqrt{m\,n}\,\frac{\|A\|_{2}}{2(2^{b-1}-1)}.$ In approximate form for large $b$, $\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
@@ B.2 Spectral Norm Error Bound
B.3 Frobenius Norm Error Bound
Theorem 4 (Frobenius Norm Error Bound for Uniform Quantization) Under the same setup as Theorem B.2, the Frobenius norm of the quantization error satisfies $\begin{split}\|A-\widehat{A}\|_{F}&\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{M}{2(2^{b-1}-1)}}\\ &\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{\|A\|_{F}}{2(2^{b-1}-1)}}.\end{split}$ In approximate form, $\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
@@ B.3 Frobenius Norm Error Bound
$\|A-\widehat{A}\|_{2}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2},$
@@ B.4 Implications for KV Cache Quantization
Spectral-Norm Perspective. Standard quantization error bounds yield
$\|\mathbf{E}_{V}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{2},$
@@
Frobenius Norm (MSE) Perspective. The Frobenius norm of the quantization error corresponds directly to the mean-squared error (MSE) when normalized by the number of elements. Specifically, for a matrix $\mathbf{A}$ with quantized approximation $\widehat{\mathbf{A}}$, the MSE is
$\mathrm{MSE}(\mathbf{A},\widehat{\mathbf{A}})\;=\;\frac{1}{\mathrm{nnz}(\mathbf{A})}\,\|\mathbf{A}-\widehat{\mathbf{A}}\|_{F}^{2},$
where $\mathrm{nnz}(\mathbf{A})$ denotes the number of entries. Thus, controlling the Frobenius norm is equivalent to controlling the MSE up to a scaling factor.
The quantization error bounds under the Frobenius norm are given by
$\|\mathbf{E}_{V}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F},$
$\|\mathbf{E}_{K}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F},$
where $L$ is the sequence length and $d_{\text{head}}$ is the head dimension.
To ensure comparable Frobenius (or MSE) errors between keys and values, we require
@@
which simplifies to
$2^{b_{V}}\,\|\mathbf{V}\|_{F}\;\approx\;2^{b_{K}}\,\|\mathbf{K}\|_{F},$
and therefore
@@
Since typically $\|\mathbf{V}\|_{F}<\|\mathbf{K}\|_{F}$, it follows that $b_{K}>b_{V}$. This reinforces the earlier spectral-norm result: keys should be allocated more bits than values to achieve balanced quantization error under an MSE criterion.
Appendix C Supplemental Results
C.1 Focus on Generative Tasks
We validate and operationalize our theorems across a diverse set of datasets, including C4, MMLU, GSM8K, EQ-Bench, CoQA, and LongBench. Our evaluation purposefully focuses on generative tasks, i.e., open-ended generation and free-form responses, because KV-cache quantization primarily affects the decoding phase rather than the prefill phase. To assess quantization error in isolation, we use C4 under a free-form decoding setup that mirrors pretraining usage, enabling us to analyze how compression directly distorts activations. To quantify downstream impact, we select GSM8K, CoQA, EQ-Bench, and LongBench, all of which rely on generate_until-style evaluation, where the model must produce extended, structured responses. In contrast, commonsense reasoning benchmarks (e.g., BoolQ, PIQA, HellaSwag, or DoRA) use log-likelihood scoring over candidate options and do not engage KV-cache quantization unless the input prompt itself is compressed. Consequently, such discriminative evaluations are orthogonal to our focus and were intentionally excluded.
Within LongBench, we surveyed all subtasks (gov_report, lcc, lsht, multi_news, narrativeqa, qasper, qmsum, repobench-p, samsum) and found that only gov_report and qmsum require substantial generation, with gov_report involving the longest outputs. Throughout this work, “LongBench” refers specifically to gov_report.
C.2 Key-Value Weight Norm
Figure 5 presents the Frobenius norms of key and value weights for the Mistral family, exhibiting the same pattern (keys consistently having higher norms than values) as shown in Figure 2 for the Llama family, further corroborating the Key-Value Norm Disparity theorem established in Section 3.1.
@@ C.2 Key-Value Weight Norm
Figure 5: Frobenius norm plot for Mistral Family. The x-axis represents the layer index in the model, while the y-axis represents the Frobenius norm magnitude. The spectral norms are higher for the key weights than for the value weights across layers.
C.3 Singular Value Distributions
@@
Figure 6 illustrates the full-spectrum singular value distribution of key and value caches across layers of the Llama 3.3 70B model on the C4 dataset. The horizontal axis indexes singular values in descending order, starting from the largest (i.e., the spectral norm), while the vertical axis shows their magnitudes. The shaded region represents the minimum-maximum range of singular values across attention heads within each layer, and the solid curves indicate the mean singular value at each rank.
@@
Figure 6: Complete singular value distribution of key and value cache for Llama 3.3-70B on the C4 dataset. The x-axis denotes the singular value indices, ordered from the largest (spectral norm) to the smallest, while the y-axis represents the corresponding magnitudes. The shaded region illustrates the range between the minimum and maximum singular values across attention heads within each layer, and the solid lines indicate the mean singular value magnitude at each index. This full-spectrum view highlights that key matrices consistently maintain significantly higher singular values throughout the entire distribution, further reinforcing their dominant representational capacity compared to value matrices.
(a) Frobenius Norm
(b) Spectral Norm
@@
Figure 7: Frobenius and spectral norms of key and value caches across layers for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the layer index. The shaded regions indicate the min-max range across attention heads within each layer, while the solid curves represent the mean norm per layer. In both cases, key matrices consistently exhibit substantially higher norms than value matrices, reflecting their stronger representational capacity.
C.4 Quantization Error
By using quantization error, we refer to evaluating how closely the quantized-then-dequantized key/value (KV) caches reconstruct their full-precision counterparts, measured as the reconstruction mean-squared error (MSE), which is exactly the Frobenius norm error normalized by the number of elements. This serves as a practical proxy for downstream quality because uniform $b$-bit quantization admits norm-based bounds in which reconstruction error scales with the cache norm and decays roughly like $2^{-b}$; thus, at a fixed bit budget, higher-norm caches incur larger distortion, and bit allocations that minimize reconstruction error are directly targeting the quantity most affected by quantization.
@@ C.4 Quantization Error
Table 4: Quantization error (mean ± std) for the key and value caches ($\mathbf{K}_{i}$ and $\mathbf{V}_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the MMLU dataset.
$\mathbf{K}_{2}$ $\mathbf{V}_{2}$ $\mathbf{K}_{3}$ $\mathbf{V}_{3}$ $\mathbf{K}_{4}$ $\mathbf{V}_{4}$
@@
Table 5: Quantization error (mean ± std) for the key and value caches ($\mathbf{K}_{i}$ and $\mathbf{V}_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the C4 dataset.
$\mathbf{K}_{2}$ $\mathbf{V}_{2}$ $\mathbf{K}_{3}$ $\mathbf{V}_{3}$ $\mathbf{K}_{4}$ $\mathbf{V}_{4}$
@@
$\mathbf{K}_{2}$ $\mathbf{V}_{2}$ $\mathbf{K}_{3}$ $\mathbf{V}_{3}$ $\mathbf{K}_{4}$ $\mathbf{V}_{4}$
@@
Figure 8: MSE of KV quantization for Llama 3.3-70B on the C4 dataset. We use quantization bit-widths ranging from 2 to 8. The x-axis represents the quantization bit-width, while the y-axis shows the MSE on a logarithmic scale. A logarithmic scale is used to highlight differences at higher bit-widths, where MSE values decrease significantly and approach zero, particularly at 8-bit. Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
(a) C4
(b) MMLU
(c) GSM8k
Figure 9: Quantization error (MSE) of key and value caches in the Llama 3.1 8B model across 32 layers for (a) C4, (b) MMLU, and (c) GSM8K. The top row shows 2-bit quantization ($\mathrm{K}_{2}$, $\mathrm{V}_{2}$), and the bottom row shows 4-bit quantization ($\mathrm{K}_{4}$, $\mathrm{V}_{4}$). Each point corresponds to an attention head within the respective layer. The x-axis denotes the layer index, and the y-axis indicates the quantization error (MSE).
@@
C.5 Downstream Accuracy Across Quantization Precisions
@@
Comprehensive downstream accuracy results for KV cache quantization using the HQQ backend are provided for CoQA (F1 scores), EQ-Parseable (parseable-pass rates and exact-match scores), and GSM8K (flexible and strict accuracy) in Tables 7,  8,  9, and 10, respectively.
[…]
Qwen3 Llama-3.2 Llama-3.2 Qwen3 Llama-3.1 Qwen3 Qwen3
0.6B 1B 3B 4B 8B 8B 32B
$\mathbf{K}_{1}\mathbf{V}_{1}$ 0.130 0.148 0.128 0.222 0.127 0.359 0.231
$\mathbf{K}_{1}\mathbf{V}_{2}$ 0.227 0.236 0.223 0.361 0.207 0.384 0.328
$\mathbf{K}_{1}\mathbf{V}_{4}$ 0.226 0.237 0.229 0.376 0.242 0.386 0.349
$\mathbf{K}_{1}\mathbf{V}_{6}$ 0.215 0.243 0.230 0.379 0.221 0.379 0.352
$\mathbf{K}_{1}\mathbf{V}_{8}$ 0.220 0.240 0.222 0.364 0.242 0.382 0.355
$\mathbf{K}_{2}\mathbf{V}_{1}$ 0.253 0.190 0.577 0.464 0.565 0.599 0.732
$\mathbf{K}_{2}\mathbf{V}_{2}$ 0.334 0.526 0.760 0.721 0.757 0.767 0.809
$\mathbf{K}_{2}\mathbf{V}_{4}$ 0.333 0.559 0.766 0.732 0.763 0.761 0.809
$\mathbf{K}_{2}\mathbf{V}_{6}$ 0.340 0.565 0.764 0.730 0.764 0.766 0.807
$\mathbf{K}_{2}\mathbf{V}_{8}$ 0.332 0.568 0.764 0.733 0.766 0.770 0.804
$\mathbf{K}_{4}\mathbf{V}_{1}$ 0.497 0.575 0.769 0.803 0.773 0.815 0.818
$\mathbf{K}_{4}\mathbf{V}_{2}$ 0.666 0.693 0.797 0.806 0.786 0.822 0.820
$\mathbf{K}_{4}\mathbf{V}_{4}$ 0.692 0.701 0.798 0.807 0.788 0.819 0.825
$\mathbf{K}_{4}\mathbf{V}_{6}$ 0.688 0.702 0.797 0.809 0.784 0.817 0.824
$\mathbf{K}_{4}\mathbf{V}_{8}$ 0.687 0.700 0.797 0.807 0.782 0.818 0.823
$\mathbf{K}_{6}\mathbf{V}_{1}$ 0.650 0.597 0.771 0.801 0.765 0.813 0.817
$\mathbf{K}_{6}\mathbf{V}_{2}$ 0.693 0.700 0.798 0.803 0.785 0.821 0.826
$\mathbf{K}_{6}\mathbf{V}_{4}$ 0.702 0.705 0.796 0.807 0.786 0.821 0.825
$\mathbf{K}_{6}\mathbf{V}_{6}$ 0.701 0.706 0.795 0.808 0.790 0.819 0.825
$\mathbf{K}_{6}\mathbf{V}_{8}$ 0.694 0.703 0.796 0.807 0.789 0.819 0.825
$\mathbf{K}_{8}\mathbf{V}_{1}$ 0.651 0.599 0.769 0.802 0.767 0.815 0.818
$\mathbf{K}_{8}\mathbf{V}_{2}$ 0.691 0.696 0.798 0.803 0.788 0.823 0.825
$\mathbf{K}_{8}\mathbf{V}_{4}$ 0.701 0.704 0.795 0.807 0.789 0.821 0.826
$\mathbf{K}_{8}\mathbf{V}_{6}$ 0.699 0.705 0.795 0.809 0.787 0.819 0.826
$\mathbf{K}_{8}\mathbf{V}_{8}$ 0.697 0.704 0.795 0.808 0.786 0.821 0.827
BF16 0.708 0.701 0.795 0.808 0.785 0.820 0.826
[…]
Qwen3 Llama-3.2 Llama-3.2 Qwen3 Llama-3.1 Qwen3 Qwen3
0.6B 1B 3B 4B 8B 8B 32B
$\mathbf{K}_{1}\mathbf{V}_{1}$ 0 0 0 0 0 0 0
$\mathbf{K}_{1}\mathbf{V}_{2}$ 0 0 0 0 0 0 0
$\mathbf{K}_{1}\mathbf{V}_{4}$ 0 0 0 0 0 0 0
$\mathbf{K}_{1}\mathbf{V}_{6}$ 0 0 0 0 0 0 0
$\mathbf{K}_{1}\mathbf{V}_{8}$ 0 0 0 0 0 0 0
$\mathbf{K}_{2}\mathbf{V}_{1}$ 0 0 8.187 0 35.673 0 2.339
$\mathbf{K}_{2}\mathbf{V}_{2}$ 0 16.374 80.702 27.485 85.965 65.497 67.836
$\mathbf{K}_{2}\mathbf{V}_{4}$ 0 34.503 84.795 39.766 88.889 70.175 70.175
$\mathbf{K}_{2}\mathbf{V}_{6}$ 0 39.181 83.041 36.842 90.059 67.836 69.591
$\mathbf{K}_{2}\mathbf{V}_{8}$ 0 38.012 84.211 40.936 88.889 68.421 70.760
$\mathbf{K}_{4}\mathbf{V}_{1}$ 0 47.953 58.480 77.778 83.626 98.830 77.778
$\mathbf{K}_{4}\mathbf{V}_{2}$ 63.743 97.076 95.322 87.719 97.661 95.322 80.702
$\mathbf{K}_{4}\mathbf{V}_{4}$ 65.497 97.076 98.830 84.795 99.415 95.322 80.702
$\mathbf{K}_{4}\mathbf{V}_{6}$ 64.328 97.076 98.830 85.380 98.830 93.567 80.702
$\mathbf{K}_{4}\mathbf{V}_{8}$ 62.573 97.076 98.830 84.795 98.830 93.567 80.702
$\mathbf{K}_{6}\mathbf{V}_{1}$ 23.392 60.234 53.801 81.287 90.643 96.491 79.532
$\mathbf{K}_{6}\mathbf{V}_{2}$ 99.415 97.076 94.737 87.719 97.661 94.737 80.702
$\mathbf{K}_{6}\mathbf{V}_{4}$ 99.415 97.661 99.415 87.135 98.830 95.906 80.702
$\mathbf{K}_{6}\mathbf{V}_{6}$ 99.415 97.661 98.830 86.550 98.830 96.491 80.702
$\mathbf{K}_{6}\mathbf{V}_{8}$ 99.415 97.661 98.830 87.135 98.830 97.661 80.702
$\mathbf{K}_{8}\mathbf{V}_{1}$ 26.901 60.234 56.725 78.947 92.983 97.076 77.193
$\mathbf{K}_{8}\mathbf{V}_{2}$ 99.415 97.076 94.737 87.135 97.661 96.491 79.532
$\mathbf{K}_{8}\mathbf{V}_{4}$ 99.415 97.661 99.415 87.719 98.830 95.322 80.702
$\mathbf{K}_{8}\mathbf{V}_{6}$ 99.415 97.661 99.415 86.550 98.830 95.906 80.702
$\mathbf{K}_{8}\mathbf{V}_{8}$ 99.415 97.661 99.415 86.550 98.830 95.906 80.702